Output (Cases) Fixed Cost Variable Cost Total cost ATC AVC MC (ДтсдQ)
0 20 - 20
1 20 12 32 32.00 12.00 12
2 20 20 40 20.00 10.00 8
3 20 28 48 16.00 9.33 8
4 20 37 57 14.25 9.25 9
5 20 47 67 13.40 9.40 10
6 20 61 81 13.50 10.17 14
7 20 81 101 14.43 11.57 20
8 20 116 136 17.00 14.50 35
9 20 171 191 21.22 19.00 55
Output(Cases) MC TC MR (ATRAQ) TR (P = $20) Total Profit (TR-TC) Marginal Profit
0 20 0 -20
1 12 32 20 20 -12 8
2 8 40 20 40 0 12
3 8 48 20 60 12 12
4 9 57 20 80 23 11
5 10 67 20 100 33 10
6 14 81 20 120 39 6
7 20 101 20 140 39 0
8 35 136 20 160 24 -15
9 55 191 20 180 -11 -35
  1. What do you think would happen if Price dropped to $15? Describe.
  2. What do you think would happen if Price increased to $25? Describe
Answer #1

If price is dropped to $15,

In a competitive market, a competitive firm increases output as long as Price is higher than or equal to marginal cost to maximize the profit.

We can see MC is lower than $15 for 6th unit of output but MC is higher than price for 7th unit of output. So, optimal output will now be 6 units (cases).

Total Cost for output of 6 units=TC=$81

Total Revenue for output of 6 units=TR=Price*output=15*6=$90

Profit=TR-TC=90-81=$9

If price is increased to $25,

In a competitive market, a competitive firm increases output as long as Price is higher than or equal to marginal cost to maximize the profit.

We can see MC is lower than $25 for 7th unit of output but MC is higher than price for 8th unit of output. So, optimal output will now be 7 units (cases).

Total Cost for output of 6 units=TC=$101

Total Revenue for output of 6 units=TR=Price*output=25*7=$175

Profit=TR-TC=175-101=$74